### v1

# Quantum Mycielskians: symmetries, twin vertices and distinguishing labelings

Abstract.
We study the properties of a generalization of the Mycielski construction to quantum graphs. We discuss quantum symmetries of quantum Mycielskians and their relation to quantum symmetries of the original quantum graphs. We introduce a quantum generalization of the notion of graphs with twin vertices and check that this concept coincides with its classical version for classical graphs with few vertices. Finally, we propose the quantum generalization of the distinguishing number and study its behavior under the quantum Mycielski construction.
Key words and phrases:
Quantum graph, Mycielskian, quantum group, twin vertices, distinguishing number
2020 Mathematics Subject Classification:
Primary: 46L89, secondary: 05C15, 46L67
1. Introduction
Classical graph theory has a variety of intriguing applications beyond pure mathematics. Its use in communication theory leads to concepts that can subsequently be generalized into the framework of quantum information theory (see, e.g. [11]). This motivated the notion of quantum graphs defined in terms of Kraus operators associated with quantum communication channels. Several aspects of quantum graphs and their characteristics have recently been widely studied; see, e.g. [6, 7, 11, 13]. Some of the concepts known from the classical graph theory that were studied in the quantum framework are chromatic numbers and clique numbers. The (classical) chromatic number is defined as the minimal number of labels (colors) that can be associated with vertices of a given graph so that the endpoints of any edge do not share the same labeling. The clique number is the maximal size of the complete graph that can be embedded as a subgraph in the original graph. For example, graphs whose clique number is equal to $2$ are triangle-free.
One of the problems in classical graph theory was to determine if it is possible to construct triangle-free graphs with an arbitrarily large chromatic number. In [16] Mycielski has shown that out of a given triangle-free graph $G$, one can construct the new one, denoted by $\mu(G)$, which is also triangle-free, contains the original one as a subgraph, and whose chromatic number is larger by one than the chromatic number of the original graph.
The natural question of extension of this result into the quantum framework has been addressed in [5], where two of the present authors proposed a construction for quantum graphs that mimics the classical construction of Mycielski. Starting from a (finite) quantum graph one can construct a new one, dubbed the quantum Mycielskian, in which the original one can be embedded. The behavior of graph characteristics under this transformation was studied in detail. While the clique number is an invariant of this operation, the (quantum) chromatic numbers satisfy only a weaker property: they can be enlarged by at most one.
In this paper, we study the further properties of this construction in the quantum framework. We begin with Section 1.1 summarizing the classical Mycielski construction. We continue in Section 1.2 with the notation for quantum graphs and the actions of quantum groups on them that will be of great use in the remaining parts of the paper. We will be especially interested in actions that preserve partitions of unity, which in the non-commutative framework correspond to colorings or labelings (see Section 4). This concept is briefly discussed in Section 1.3. At the end of Section 1 we recall the construction of quantum Mycielskian. In Section 2 we discuss isomorphisms of quantum Mycielskians, while Section 3 is devoted to the study of how the Mycielski construction for quantum graphs affects their symmetries. This leads us to the concept of quantum twin vertices that we then study in Sections 3.1 and 3.2 (as well as in the Appendix). Yet another characteristic of classical graphs that can also be defined in the quantum setting is that of distinguishing labeling. We define the distinguishing number in Section  4 and study the impact of the quantum Mycielski construction.
1.1. Classical Mycielski construction
Let $G=(V_{G},E_{G})$ be a (finite, undirected) graph. The Mycielskian of $G$, denoted by $\mu(G)$ is a graph with the set of vertices $V_{\mu(G)}=\{\bullet\}\sqcup V_{G}\sqcup V_{G}$, where $\bullet$ denotes a distinguished point, called master vertex. The edge set $E_{\mu(G)}$ can be specified by the adjacency matrix $A_{\mu(G)}$. We recall that the adjacency matrix $A_{G}$ for an undirected graph $G$ is a $\{0,1\}$-matrix whose columns and rows are indexed by the set of vertices, and $(A_{G})_{i,j}$ is non-zero if and only if $\{v_{i},v_{j}\}\in E$. For the Mycielskian, we have
$A_{\mu(G)}=\begin{bmatrix}0&\vec{0}^{T}&\vec{1}^{T}\\ \vec{0}&A_{G}&A_{G}\\ \vec{1}&A_{G}&0\end{bmatrix},$
where $\vec{1}$ (resp. $\vec{0}$) is the column vector with all entries being $1$ (resp. 00). Mycielski’s theorem [16] says that denoting by $\omega(\cdot)$ the clique number of a graph we have $\omega(\mu(G))=\omega(G)$, i.e. the clique number remains unchanged under the passage from $G$ to $\mu(G)$, and for the chromatic numbers $\chi$ we have $\chi(\mu(G))=\chi(G)+1$.
The Mycielski construction can be generalized to so-called $r$-Mycielskian ($r\geq 2$) also known as higher cones over graphs ([20, 21]). In this case,
$V_{\mu_{r-1}(G)}=\{\bullet\}\sqcup V^{1}\sqcup\ldots\sqcup V^{n}$
with $V^{j}=\{v_{1}^{j},\ldots,v^{j}_{|V_{G}|}\}$, $j=1,\ldots,r$ denoting distinct copies of $V_{G}$, and
$E=E_{G}\cup\bigcup_{i=0}^{r-1}\bigl{\{}\{v^{i}_{j},v^{i+1}_{j^{\prime}}\}\,\bigr{|}\bigl{.}\,\{v^{0}_{j},v^{0}_{j^{\prime}}\}\in E_{G}\bigr{\}}\cup\bigl{\{}\{v^{r}_{1},\bullet\},\{v^{r}_{2},\bullet\},\ldots,\{v^{r}_{|V_{G}|},\bullet\}\bigr{\}}.$
1.2. Quantum graphs and their quantum symmetries
Definition 1.1.
Let $\mathsf{A}$ be a finite-dimensional $\mathrm{C}^{*}$-algebra. A $\delta$-form on $\mathsf{A}$ is a faithful state $\psi$ on $\mathsf{A}$ such that writing $L^{2}(\mathsf{A},\psi)$ for the GNS Hilbert space for $\psi$ and regarding the multiplication map $m\colon\mathsf{A}\otimes\mathsf{A}\to\mathsf{A}$ as a linear operator $L^{2}(\mathsf{A},\psi)\otimes L^{2}(\mathsf{A},\psi)\to L^{2}(\mathsf{A},\psi)$ we have
$mm^{*}=\delta(\psi)^{2}\mathrm{id}_{L^{2}(\mathsf{A},\psi)}$
for some positive constant $\delta(\psi)$. Since $\mathsf{A}$ and $L^{2}(\mathsf{A},\psi)$ are naturally isomorphic as vector spaces, we usually write $\mathrm{id}_{\mathsf{A}}$ instead of $\mathrm{id}_{L^{2}(\mathsf{A},\psi)}$ even when viewed as a map of Hilbert spaces.
A quantum graph $\mathcal{G}$ is specified by elements of a triple $(\mathbb{V}_{\scriptscriptstyle\mathcal{G}},\psi_{\scriptscriptstyle\mathcal{G}},A_{\scriptscriptstyle\mathcal{G}})$, where
  • $\mathbb{V}_{\scriptscriptstyle\mathcal{G}}$ is a finite quantum space,
  • $\psi_{\scriptscriptstyle\mathcal{G}}$ is a $\delta$-form on the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})$,
  • $A_{\scriptscriptstyle\mathcal{G}}$ is a self-adjoint operator on the GNS Hilbert space $L^{2}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ such that
    $m(A\otimes A)m^{*}=A\quad\text{and}\quad A=(\mathrm{id}_{\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})}\otimes\eta^{*}m)(\mathrm{id}_{\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})}\otimes A\otimes\mathrm{id}_{\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})})(m^{*}\eta\otimes\mathrm{id}_{\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})}),$
    where
    $m\colon\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})\otimes\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})\longrightarrow\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})\quad\text{and}\quad\eta\colon\mathbb{C}\longrightarrow\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})$
    are the multiplication and unit maps regarded as linear operators between Hilbert spaces $L^{2}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})\otimes L^{2}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})\to L^{2}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ and $\mathbb{C}\to L^{2}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ respectively.
In what follows, the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})$ will be denoted by $\operatorname{C}(\mathcal{G})$ and the Hilbert space $L^{2}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ by $L^{2}(\mathcal{G})$. We will also shorten $\delta(\psi_{\scriptscriptstyle\mathcal{G}})$ to $\delta_{\scriptscriptstyle\mathcal{G}}$.
Definition 1.2 ([7, Definition 6.3]).
Let $\mathcal{G}$ be a quantum graph. The quantum symmetry group of $\mathcal{G}$ is the compact quantum group $\operatorname{QAut}(\mathcal{G})$ equipped with a $\psi_{\scriptscriptstyle\mathcal{G}}$-preserving action $\rho_{\scriptscriptstyle\mathcal{G}}\colon\operatorname{C}(\mathcal{G})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ such that $\rho_{\scriptscriptstyle\mathcal{G}}\circ{A_{\scriptscriptstyle\mathcal{G}}}=(A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\circ\rho_{\scriptscriptstyle\mathcal{G}}$ and which is universal among all $\psi_{\scriptscriptstyle\mathcal{G}}$-preserving compact quantum group actions $\alpha\colon\operatorname{C}(\mathcal{G})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{C}(\mathbb{G})$ such that $\alpha\circ A_{\scriptscriptstyle\mathcal{G}}=(A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\mathbb{G})})\circ\alpha$.
Any choice of a basis $\{e_{1},\dotsc,e_{n}\}$ in $\operatorname{C}(\mathcal{G})$ determines elements $u_{i,j}$ of $\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ by
$\rho_{\scriptscriptstyle\mathcal{G}}(e_{j})=\sum_{i=1}^{n}e_{i}\otimes u_{i,j}.$
If the basis $\{e_{1},\dotsc,e_{n}\}$ is orthonormal in $L^{2}(\mathcal{G})$ then the matrix $u$ with elements $u_{i,j}$ is unitary. In all cases it is a finite-dimensional representation of $\operatorname{QAut}(\mathcal{G})$ and $\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ is generated by $\{u_{i,j}\,|\,i,j=1,\dotsc,\dim{\operatorname{C}(\mathcal{G})}\}$. The compatibility of $\rho_{\scriptscriptstyle\mathcal{G}}$ with $A_{\scriptscriptstyle\mathcal{G}}$ is equivalent to $(\mathds{1}\otimes{A_{\scriptscriptstyle\mathcal{G}}})u=u(\mathds{1}\otimes{A_{\scriptscriptstyle\mathcal{G}}})$.
The quantum symmetry group of a quantum graph $\mathcal{G}$ is by construction a quantum subgroup of a well-studied quantum group $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ of symmetries of the quantum space $\mathbb{V}_{\scriptscriptstyle\mathcal{G}}$ together with the $\delta$-form $\psi_{\scriptscriptstyle\mathcal{G}}$ (this larger quantum group does not preserve the adjacency matrix). The quantum groups of symmetries of finite quantum spaces equipped with a $\delta$-form were introduced in [2].
1.3. Actions preserving a partition of unity
We now briefly introduce a construction not directly related to quantum graphs which will prove useful later.
Let $\mathsf{A}$ be a unital $\mathrm{C}^{*}$-algebra and let $\rho\colon\mathsf{A}\to\mathsf{A}\otimes\operatorname{C}(\mathbb{G})$ be an action of a compact quantum group $\mathbb{G}$ on the quantum space underlying $\mathsf{A}$. Let $\mathfrak{P}=\{P_{1},\dotsc,P_{n}\}$ be a partition of unity in $\mathsf{A}$. Let $\mathsf{J}$ be the ideal in $\operatorname{C}(\mathbb{G})$ generated by the elements
$\bigl{\{}(\omega\otimes\mathrm{id})\rho(P_{a})-\omega(P_{a})\mathds{1}\,\bigr{|}\bigl{.}\,\omega\in\mathsf{A}^{*},\>a=1,\dotsc,n\bigr{\}}.$
Denote the quotient $\operatorname{C}(\mathbb{G})/\mathsf{J}$ by $\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})$ and let $\pi\colon\operatorname{C}(\mathbb{G})\to\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})$ be the corresponding quotient map. Furthermore, let $\rho_{\scriptscriptstyle{\mathfrak{P}}}=(\mathrm{id}\otimes\pi)\circ\rho$. Then $\rho_{\scriptscriptstyle{\mathfrak{P}}}\colon\mathsf{A}\to\mathsf{A}\otimes\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})$ and $\rho_{\scriptscriptstyle{\mathfrak{P}}}(P_{a})=P_{a}\otimes\mathds{1}$ for all $a$ (since slice maps $(\omega\otimes\mathrm{id})$ separate elements of $\mathsf{A}\otimes\operatorname{C}(\mathbb{G})$).
Next, let $\mathbb{Y}$ be any compact quantum space and let $\varphi\colon\operatorname{C}(\mathbb{G})\to\operatorname{C}(\mathbb{Y})$ be a unital *-homomorphism such that $(\mathrm{id}\otimes\varphi)\rho(P_{a})=P_{a}\otimes\mathds{1}$ for all $a$. Then for any $\omega\in\mathsf{A}^{*}$ and any $a$ we have
$0=(\omega\otimes\mathrm{id})\bigl{(}(\mathrm{id}\otimes\varphi)\rho(P_{a})-P_{a}\otimes\mathds{1}\bigr{)}=\varphi\bigl{(}(\omega\otimes\mathrm{id})\rho(P_{a})-\omega(P_{a})\mathds{1}\bigr{)},$
so $\mathsf{J}\subset\ker\varphi$. It follows that there exists a unique $\Lambda\colon\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})\to\operatorname{C}(\mathbb{Y})$ such that $\varphi=\Lambda\circ\pi$.
Now take $\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}\times\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}$ for $\mathbb{Y}$ in the sense that $\operatorname{C}(\mathbb{Y})=\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})\otimes\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})$ and let $\varphi=(\pi\otimes\pi)\circ\Delta_{\mathbb{G}}$, where $\Delta_{\mathbb{G}}$ is the comultiplication on $\operatorname{C}(\mathbb{G})$. Then by the reasoning above there exists a unique $\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}}\colon\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})\to\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})\otimes\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}})$ such that $(\pi\otimes\pi)\circ\Delta_{\mathbb{G}}=\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}}\circ\pi$. Moreover, we have
$(\rho_{\scriptscriptstyle{\mathfrak{P}}}\otimes\mathrm{id})\circ\rho_{\scriptscriptstyle{\mathfrak{P}}}=(\mathrm{id}\otimes\varphi)\circ\rho=(\mathrm{id}\otimes\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}})\circ\rho_{\scriptscriptstyle{\mathfrak{P}}}.$
Of course
[…]
so $\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}}$ is coassociative, since $\pi$ is surjective. Finally, applying $\pi\otimes\pi$ to both sides of the equalities
$\overline{\bigl{(}\operatorname{C}(\mathbb{G})\otimes\mathds{1}\bigr{)}\Delta_{\mathbb{G}}\bigl{(}\operatorname{C}(\mathbb{G})\bigr{)}}=\operatorname{C}(\mathbb{G})\otimes\operatorname{C}(\mathbb{G})\quad\text{and}\quad\overline{\Delta_{\mathbb{G}}\bigl{(}\operatorname{C}(\mathbb{G})\bigr{)}\bigl{(}\mathds{1}\otimes\operatorname{C}(\mathbb{G})\bigr{)}}=\operatorname{C}(\mathbb{G})\otimes\operatorname{C}(\mathbb{G})$
we find that the density conditions are satisfied for $(\operatorname{C}(\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}),\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}})$ and consequently $\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}$ is a compact quantum group.
1.4. Mycielskians of quantum graphs
Here we recall the construction introduced in [5]. For $r\geq{2}$ the Mycielski construction of order $r-1$ for quantum graphs is the procedure of passing from a quantum graph $\mathcal{G}$ to a new quantum graph $\mu_{r-1}(\mathcal{G})$ defined as follows ([5, Section 4]):
  • the quantum space $\mathbb{V}_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}$ is the disjoint union of a point and $r$ copies of $\mathbb{V}_{\scriptscriptstyle\mathcal{G}}$: $\operatorname{C}(\mu_{r-1}(\mathcal{G}))=\mathbb{C}\oplus\bigoplus\limits_{k=1}^{r}\operatorname{C}(\mathcal{G})$,
  • the state $\psi_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}$ is defined by the formula
    $\psi_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\left(\left[\begin{smallmatrix}\lambda\\ x_{1}\\[-2.0pt] \vdots\\[2.0pt] x_{r}\end{smallmatrix}\right]\right)=\tfrac{1}{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}\biggl{(}\lambda+\delta_{\scriptscriptstyle\mathcal{G}}^{2}\sum_{k=1}^{r}\psi_{\scriptscriptstyle\mathcal{G}}(x_{i})\biggr{)}$
    (we note the obvious identification $L^{2}(\mu_{r-1}(\mathcal{G}))=\mathbb{C}\oplus\bigoplus\limits_{k=1}^{r}L^{2}(\mathcal{G})$ as vector spaces),
  • the quantum adjacency matrix $A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}$ acts on $L^{2}(\mu_{r-1}(\mathcal{G}))$ as
    $A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\left[\begin{smallmatrix}\lambda\\ x_{1}\\[-2.0pt] \vdots\\[2.0pt] x_{r}\end{smallmatrix}\right]=\left[\begin{smallmatrix}\delta_{\scriptscriptstyle\mathcal{G}}^{2}\psi_{\scriptscriptstyle\mathcal{G}}(x^{r})\\ A_{\scriptscriptstyle\mathcal{G}}(x_{1}+x_{2})\\ A_{\scriptscriptstyle\mathcal{G}}(x_{2}+x_{3})\\[-2.0pt] \vdots\\[2.0pt] A_{\scriptscriptstyle\mathcal{G}}(x_{r-2}+x_{r-1})\\ \lambda\mathds{1}_{\operatorname{C}(\mathcal{G})}+A_{\scriptscriptstyle\mathcal{G}}x_{r-1}\end{smallmatrix}\right].$
Additionally, for $r=1$ we let $\mu_{r-1}(\mathcal{G})=\mu_{0}(\mathcal{G})$ be the original quantum graph $\mathcal{G}$. The quantum graph $\mu_{r-1}(\mathcal{G})$ contains $r$ “copies” of the original graph $\mathcal{G}$. In particular $\mu_{1}(\mathcal{G})$ also denoted $\mu(\mathcal{G})$ is a non-commutative generalization of the Mycielskian of a finite graph ([16]).
We let $\iota_{0}\colon\mathbb{C}\to L^{2}(\mu_{r-1}(\mathcal{G}))$ and $\iota_{1},\dotsc,\iota_{r}\colon L^{2}(\mathcal{G})\to L^{2}(\mu_{r-1}(\mathcal{G}))$ denote the isometric embeddings onto the respective direct summands and $\iota_{0}^{*}\colon L^{2}(\mu_{r-1}(\mathcal{G}))\to\mathbb{C}$ and $\iota_{1}^{*},\dotsc,\iota_{r}^{*}\colon L^{2}(\mu_{r-1}(\mathcal{G}))\to L^{2}(\mathcal{G})$ are their hermitian adjoints:
$\iota_{0}(\lambda)$ $=\sqrt{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}\left[\begin{smallmatrix}\lambda\\ 0\\[-5.0pt] \vdots\\ 0\end{smallmatrix}\right],$ $\iota_{0}^{*}\left[\begin{smallmatrix}\lambda\\ x_{1}\\[-2.0pt] \vdots\\[2.0pt] x_{r}\end{smallmatrix}\right]$ $=\sqrt{\tfrac{1}{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\,\lambda,$
$\iota_{k}(x)$ $=\sqrt{\tfrac{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}{\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\left[\begin{smallmatrix}0\\[-5.0pt] \vdots\\ x\\[-5.0pt] \vdots\\ 0\end{smallmatrix}\right],$ $\iota_{k}^{*}\left[\begin{smallmatrix}\lambda\\ x_{1}\\[-2.0pt] \vdots\\[2.0pt] x_{r}\end{smallmatrix}\right]$ $=\sqrt{\tfrac{\delta_{\scriptscriptstyle\mathcal{G}}^{2}}{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\,x_{k}$
for $k=1,\dotsc,r$. Note that although $\iota_{0},\dotsc,\iota_{r}$, and their adjoints are not multiplicative, they are all *-maps. In this notation, we have (cf. [5, Eq. (9)]):
$A_{\mu_{r-1}(\mathcal{G})}=\delta_{\mathcal{G}}\iota_{r}\eta_{\mathcal{G}}\iota_{0}^{\ast}+\delta_{\mathcal{G}}\iota_{0}\eta_{\mathcal{G}}^{\ast}\iota_{r}^{\ast}+\iota_{1}A_{\mathcal{G}}\iota_{1}^{\ast}+\sum\limits_{k=1}^{r-1}\bigl{(}\iota_{k}A_{\mathcal{G}}\iota_{k+1}^{\ast}+\iota_{k+1}A_{\mathcal{G}}\iota_{k}^{\ast}\bigr{)}.$
In what follows, we will freely identify the $\mathrm{C}^{*}$-algebras of functions on the quantum sets of vertices of the quantum graph with the Hilbert spaces obtained via the GNS construction for the corresponding $\delta$-forms. In particular, all linear maps between the considered $\mathrm{C}^{*}$-algebras (homomorphisms and others) may be understood as maps of Hilbert spaces, and we will have at our disposal their hermitian adjoints.
2. Quantum isomorphisms of Mycielskians
Let $\mathcal{G}$ and $\mathcal{F}$ be quantum graphs. We say that $\mathcal{G}$ and $\mathcal{F}$ are quantum isomorphic if the quantum groups $\operatorname{QAut}(\mathcal{G})$ and $\operatorname{QAut}(\mathcal{F})$ are monoidally equivalent ([6, Section 4] and [7, Definition 6.5]). The notion of monoidal equivalence of compact quantum groups is described e.g. in [10].
Proposition 2.1.
Let $\mathcal{F}$ and $\mathcal{G}$ be quantum graphs. If $\mathcal{F}$ and $\mathcal{G}$ are quantum isomorphic, then $\delta_{\mathcal{F}}=\delta_{\scriptscriptstyle\mathcal{G}}$.
Proof.
The monoidal equivalence of the quantum groups $\operatorname{QAut}(\mathcal{G})$ and $\operatorname{QAut}(\mathcal{F})$ amounts to a certain algebra $\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))$ being non-zero ([7, Theorem 6.6]). Since this algebra is a homomorphic image for the corresponding linking algebra for the quantum automorphism groups $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\mathcal{F}}),\psi_{\mathcal{F}})$ and $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ (see [10, Section 4]), the latter algebra is non-zero and hence $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\mathcal{F}}),\psi_{\mathcal{F}})$ and $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ are monoidally equivalent. Then, the result follows from [10, Theorem 4.7]. ∎
Proposition 2.2.
Let $\mathcal{G}$ and $\mathcal{F}$ be quantum isomorphic. Then $\mu_{r-1}(\mathcal{G})$ and $\mu_{r-1}(\mathcal{F})$ are quantum isomorphic for any $r\geq 1$.
Proof.
Let
$p=[p_{ij}]_{i,j}\in\operatorname{B}\bigl{(}L^{2}(\mathcal{G}),L^{2}(\mathcal{F})\bigr{)}\otimes\mathcal{O}\bigl{(}\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F})\bigr{)}$
be the unitary matrix defining $\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))$ as in [6, Definition 4.1] (remember the latter is non-zero) and let
$\rho_{\mathcal{F},\mathcal{G}}\colon\operatorname{C}(\mathcal{G})\ni e_{j}\longmapsto\sum\limits_{i}f_{i}\otimes p_{ij}\in\operatorname{C}(\mathcal{F})\otimes\mathcal{O}\bigl{(}\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F})\bigr{)}$
be the corresponding unital *-homomorphism intertwining the quantum adjacency matrices. By Proposition 2.1 we know that $\delta_{\scriptscriptstyle\mathcal{G}}=\delta_{\mathcal{F}}$ and for the remainder of the proof we will denote this value by $\delta$. The maps introduced above can be represented as
$p\colon L^{2}(\mathcal{G})\otimes\mathcal{O}\bigl{(}\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F})\bigr{)}\ni\xi\otimes a\longmapsto\sum\limits_{i,j}{\left|f_{i}\right\rangle}{\left\langle{e_{j}}\,\vline\,\xi\right\rangle}\otimes p_{ij}a\in L^{2}(\mathcal{F})\otimes\mathcal{O}\bigl{(}\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F})\bigr{)}$
and
$\rho_{\mathcal{F},\mathcal{G}}(\xi)=p(\xi\otimes 1).$
We define a map $\widehat{p}\in\operatorname{B}(L^{2}(\mu_{r-1}(\mathcal{G})),L^{2}(\mu_{r-1}(\mathcal{F})))\otimes\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))$:
$\widehat{p}=\varpi_{\mathcal{F},0}\varpi_{\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\mathcal{F},k}p\varpi_{\mathcal{G},k}^{*},$
where $\varpi_{\mathcal{G},j}=\iota_{\mathcal{G},j}\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))}$ and $\varpi_{\mathcal{F},j}=\iota_{\mathcal{F},j}\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))}$, with $\iota_{\mathcal{G},j}$ and $\iota_{\mathcal{F},j}$ denoting the isometric embeddings of $L^{2}(\mathcal{G})$ and $L^{2}(\mathcal{F})$ into $L^{2}(\mu_{r-1}(\mathcal{G}))$ and $L^{2}(\mu_{r-1}(\mathcal{F}))$, respectively.
First, notice that
$\widehat{p}\widehat{p}^{*}=\varpi_{\mathcal{F},0}\varpi_{\mathcal{F},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\mathcal{F},k}pp^{*}\varpi_{\mathcal{F},k}^{*}=\sum\limits_{j=0}^{r}\varpi_{\mathcal{F},j}\varpi_{\mathcal{F},j}^{*}=\mathrm{id}_{L^{2}(\mu_{r-1}(\mathcal{F}))},$
$\widehat{p}^{*}\widehat{p}=\varpi_{\mathcal{G},0}\varpi_{\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{k,\mathcal{G}}pp^{*}\varpi_{\mathcal{G},k}^{*}=\sum\limits_{j=0}^{r}\varpi_{\mathcal{G},j}\varpi_{\mathcal{G},j}^{*}=\mathrm{id}_{L^{2}(\mu_{r-1}(\mathcal{G}))},$
so that $\widehat{p}$ is unitary.
Next, we check the intertwining property. First, we compute
$\widehat{p}A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}$ $=\biggl{(}\varpi_{\mathcal{F},0}\varpi_{\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\mathcal{F},k}p\varpi_{\mathcal{G},k}^{*}\biggr{)}\biggl{(}\delta\varpi_{\mathcal{G},r}\eta_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},0}^{*}+\delta\varpi_{\mathcal{G},0}\eta^{*}_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},r}^{*}+\varpi_{\mathcal{G},1}A_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},1}^{*}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum\limits_{k=1}^{r-1}\bigl{(}\varpi_{\mathcal{G},k}A_{\mathcal{G}}\varpi_{\mathcal{G},k+1}^{*}+\varpi_{\mathcal{G},k+1}A_{\mathcal{G}}\varpi_{\mathcal{G},k}^{*}\bigr{)}\biggr{)}$
$=\delta\varpi_{\mathcal{F},0}\eta_{\scriptscriptstyle\mathcal{G}}^{*}\varpi_{\mathcal{G},r}^{*}+\delta\varpi_{\mathcal{F},r}p\eta_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},0}^{*}+\varpi_{\mathcal{F},1}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},1}^{*}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum\limits_{j=1}^{r-1}\bigl{(}\varpi_{j,\mathcal{F}}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},j+1}^{*}+\varpi_{\mathcal{F},j+1}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},j}^{*}\bigr{)},$
and similarly,
$A_{\mu_{r-1}(\mathcal{F})}\widehat{p}=\delta\varpi_{\mathcal{F},0}\eta_{\mathcal{F}}^{*}p\varpi_{\mathcal{G},r}^{*}+\delta\varpi_{\mathcal{F},r}\eta_{\mathcal{F}}\varpi_{\mathcal{G},0}^{*}$ $+\varpi_{\mathcal{F},1}A_{\mathcal{F}}p\varpi_{\mathcal{G},1}^{*}$
$+\sum\limits_{j=1}^{r-1}\bigl{(}\varpi_{j,\mathcal{F}}A_{\mathcal{F}}p\varpi_{\mathcal{G},j+1}^{*}+\varpi_{\mathcal{F},j+1}A_{\mathcal{F}}p\varpi_{\mathcal{G},j}^{*}\bigr{)}.$
Since $p\eta_{\scriptscriptstyle\mathcal{G}}=\eta_{\mathcal{F}}$, we also have $\eta_{\scriptscriptstyle\mathcal{G}}^{*}=\eta_{\mathcal{F}}^{*}p$. This shows that indeed the intertwining property holds. Finally, the map
$\rho_{\mu_{r-1}(\mathcal{F}),\mu_{r-1}(\mathcal{G})}\left(\left[\begin{smallmatrix}\lambda\\ \xi_{1}\\[-3.0pt] \vdots\\[1.0pt] \xi_{r}\end{smallmatrix}\right]\right)=\widehat{p}\left(\left[\begin{smallmatrix}\lambda\\ \xi_{1}\\[-3.0pt] \vdots\\[1.0pt] \xi_{r}\end{smallmatrix}\right]\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))}\right)$
is a unital *-homomorphism by construction. ∎
3. Quantum symmetries of $\mu_{r-1}(\mathcal{G})$ from quantum symmetries of $\mathcal{G}$
Our goal now is to determine how the Mycielski transformation affects the quantum symmetries of the original quantum graph. We start with the following:
Proposition 3.1.
Let $\mathcal{G}$ be a quantum graph. For $r\geq{2}$ define the linear map $\Phi_{r}\colon\operatorname{C}(\mu_{r-1}(\mathcal{G}))\to\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ by
$\Phi_{r}(x)=\bigl{(}\iota_{0}\iota_{0}^{*}(x)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\bigr{)}+\sum_{k=1}^{r}(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}\bigl{(}\iota_{k}^{*}(x)\bigr{)}.$
Then $\Phi_{r}$ is a unital *-homomorphism and denoting by $\Delta$ the comultiplication on $\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ we have
$(\Phi_{r}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\circ\Phi_{r}$ $=(\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\otimes\Delta)\circ\Phi_{r},$ (1a)
$\Phi_{r}\circ A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}$ $=(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\circ\Phi_{r},$ (1b)
Moreover the set $\bigl{\{}(\omega\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\Phi_{r}(x)\,\bigr{|}\bigl{.}\,x\in\operatorname{C}(\mu_{r-1}(\mathcal{G})),\>\omega\in\operatorname{C}(\mu_{r-1}(\mathcal{G}))^{*}\bigr{\}}$ generates the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$.
Proof.
We first establish that $\Phi_{r}$ is a unital *-homomorphism. We have
$\mathds{1}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))}=\sqrt{\tfrac{1}{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\biggl{(}\iota_{0}(1)+\delta_{\scriptscriptstyle\mathcal{G}}\sum_{k=1}^{r}\iota_{k}(\mathds{1}_{\operatorname{C}(\mathcal{G})})\biggr{)}.$
Hence
$\Phi_{r}$ $(\mathds{1}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))})=\sqrt{\tfrac{1}{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\biggl{(}\iota_{0}(1)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}+\delta_{\scriptscriptstyle\mathcal{G}}\sum_{k,l=1}^{r}(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}\bigl{(}\iota_{k}^{*}\iota_{l}(\mathds{1}_{\operatorname{C}(\mathcal{G})})\bigr{)}\biggr{)}$
$\qquad=\sqrt{\tfrac{1}{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\biggl{(}\iota_{0}(1)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}+\delta_{\scriptscriptstyle\mathcal{G}}\sum_{k=1}^{r}(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\mathds{1}_{\operatorname{C}(\mathcal{G})}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\biggr{)}$
$\qquad=\sqrt{\tfrac{1}{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\biggl{(}\iota_{0}(1)+\delta_{\scriptscriptstyle\mathcal{G}}\sum_{k=1}^{r}\iota_{k}(\mathds{1}_{\operatorname{C}(\mathcal{G})})\biggr{)}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}=\mathds{1}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}.$
To prove the multiplicativity of $\Phi_{r}$ we temporarily denote by $m^{(0)}$, $m^{(1)}$, $m^{(2)}$ and $m^{(3)}$ the multiplication maps on $\mathbb{C}$, $\operatorname{C}(\mathcal{G})$, $\operatorname{C}(\mu_{r-1}(\mathcal{G}))$ and $\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ respectively (the last one defined only on the algebraic tensor product) and recall from [5, Section 4] that
$m^{(2)}=\sqrt{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}\biggl{(}\iota_{0}m^{(0)}(\iota_{0}^{*}\otimes\iota_{0}^{*})+\delta_{\scriptscriptstyle\mathcal{G}}^{-1}\sum_{k=1}^{r}\iota_{k}m^{(1)}(\iota_{k}^{*}\otimes\iota_{k}^{*})\biggr{)}.$
Then note that for $s\in\{1,\dotsc,r\}$ the map $\sqrt{\tfrac{\delta_{\scriptscriptstyle\mathcal{G}}^{2}}{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\iota_{s}$ is multiplicative, which means that
$m^{(2)}\circ(\iota_{s}\otimes\iota_{s})=\sqrt{\tfrac{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}{\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\iota_{s}\circ m^{(1)}.$ (2)
Hence for $(p,q)\neq(0,0)$, with $s=\max\{p,q\}$ and $\boldsymbol{\chi}$ denoting the flip map, by (2) we have
[…]
Similarly we compute that $\Phi_{r}\circ m^{(2)}=(m^{(2)}\otimes m^{(3)})(\mathrm{id}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))}\otimes\boldsymbol{\chi}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\Phi_{r}\otimes\Phi_{r})$ on simple tensors of the form $\iota_{0}(\lambda)\otimes\iota_{0}(\lambda^{\prime})$ for any $\lambda,\lambda^{\prime}\in\mathbb{C}$. Since $\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{C}(\mu_{r-1}(\mathcal{G}))$ is spanned by elements of the form $\iota_{k}(a)\otimes\iota_{l}(b)$ with $k,l\in\{0,\dotsc,r\}$, these computations show that $\Phi_{r}$ is multiplicative.
The fact that $\Phi_{r}$ is a *-map follows immediately from the remark in Section 1.4 that all maps involved in the definition of $\Phi_{r}$ are *-maps.
To check (1a) we note that for $\lambda\in\mathbb{C}$ we have $\Phi_{r}(\iota_{0}(\lambda))=\iota_{0}(\lambda)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$ which immediately shows that $(\Phi_{r}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\Phi_{r}(\iota_{0}(\lambda))=(\mathrm{id}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))}\otimes\Delta)\Phi_{r}(\iota_{0}(\lambda))$ and similarly, since for any $k\in\{1,\dotsc,r\}$ we have $\Phi_{r}\circ\iota_{k}=(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\circ\rho_{\scriptscriptstyle\mathcal{G}}$, one finds that
$(\Phi_{r}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\Phi_{r}\bigl{(}\iota_{k}(a)\bigr{)}$ $=(\Phi_{r}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\rho_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\mathrm{id}_{\operatorname{C}(\mathcal{G})}\otimes\Delta)\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=(\mathrm{id}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))}\otimes\Delta)(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=(\mathrm{id}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))}\otimes\Delta)\Phi_{r}\bigl{(}\iota_{k}(a)\bigr{)}$
for any $a\in\operatorname{C}(\mathcal{G})$.
We now move on to the proof of (1b). To do this, we recall:
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\eta_{\mathcal{G}}\iota_{0}^{*}+\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}\eta_{\mathcal{G}}^{*}\iota_{r}^{*}+\iota_{1}A_{\scriptscriptstyle\mathcal{G}}\iota_{1}^{*}+\sum_{k=1}^{r-1}(\iota_{k}A_{\scriptscriptstyle\mathcal{G}}\iota_{k+1}^{*}+\iota_{k+1}A_{\scriptscriptstyle\mathcal{G}}\iota_{k}^{*}),$
where $\eta_{\mathcal{G}}$ is the unit map $\mathbb{C}\to\operatorname{C}(\mathcal{G})$. It follows from this equation that
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{0}$ $=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\eta_{\mathcal{G}},$
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{1}$ $=\iota_{1}A_{\scriptscriptstyle\mathcal{G}}+\iota_{2}A_{\scriptscriptstyle\mathcal{G}},$
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{l}$ $=\iota_{l-1}A_{\scriptscriptstyle\mathcal{G}}+\iota_{l+1}A_{\scriptscriptstyle\mathcal{G}},\qquad l=2,\dotsc,r-1,$
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{r}$ $=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}\eta_{\mathcal{G}}^{*}+\iota_{r-1}A_{\scriptscriptstyle\mathcal{G}}.$
Using this we check that for $\lambda\in\mathbb{C}$
$(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})$ $\Phi_{r}\bigl{(}\iota_{0}(\lambda)\bigr{)}=A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{0}(\lambda)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\bigl{(}\eta_{\mathcal{G}}(\lambda)\bigr{)}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$
$=\delta_{\scriptscriptstyle\mathcal{G}}\lambda\iota_{r}(\mathds{1}_{\operatorname{C}(\mathcal{G})})\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}=\delta_{\scriptscriptstyle\mathcal{G}}\lambda(\iota_{r}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(\mathds{1}_{\operatorname{C}(\mathcal{G})})$
$=\delta_{\scriptscriptstyle\mathcal{G}}\lambda\Phi_{r}\bigl{(}\iota_{r}(\mathds{1}_{\operatorname{C}(\mathcal{G})})\bigr{)}=\delta_{\scriptscriptstyle\mathcal{G}}\Phi_{r}\bigl{(}\iota_{r}(\lambda\mathds{1}_{\operatorname{C}(\mathcal{G})})\bigr{)}=\Phi_{r}\bigl{(}A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{0}(\lambda)\bigr{)}.$
Similarly for $a\in\operatorname{C}(\mathcal{G})$ we have
$(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})$ $\Phi_{r}\bigl{(}\iota_{1}(a)\bigr{)}=(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=(\iota_{1}A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)+(\iota_{2}A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=(\iota_{1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)+(\iota_{2}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)$
$=\Phi_{r}\bigl{(}\iota_{1}(A_{\scriptscriptstyle\mathcal{G}}a)+\iota_{2}(A_{\scriptscriptstyle\mathcal{G}}a)\bigr{)}=\Phi_{r}\bigl{(}A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{1}(a)\bigr{)},$
as well as
$(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})$ $\Phi_{r}\bigl{(}\iota_{l}(a)\bigr{)}=(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{l}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=(\iota_{l-1}A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)+(\iota_{l+1}A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=(\iota_{l-1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)+(\iota_{l+1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)$
$=\Phi_{r}\bigl{(}\iota_{l-1}(A_{\scriptscriptstyle\mathcal{G}}a)+\iota_{l+1}(A_{\scriptscriptstyle\mathcal{G}}a)\bigr{)}=\Phi_{r}\bigl{(}A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{l}(a)\bigr{)}$
for $l=2,\dotsc,r-1$, and finally using the fact that $\eta_{\mathcal{G}}^{*}=\psi$ and that $\psi$ is $\rho_{\scriptscriptstyle\mathcal{G}}$-invariant we arrive at
$(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\>\otimes\>$ $\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\Phi_{r}\bigl{(}\iota_{r}(a)\bigr{)}=(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{r}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=\delta_{\scriptscriptstyle\mathcal{G}}(\iota_{0}\eta_{\mathcal{G}}^{*}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)+(\iota_{r-1}A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
$=\delta_{\scriptscriptstyle\mathcal{G}}(\iota_{0}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\psi\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)+(\iota_{r-1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)$
$=\delta_{\scriptscriptstyle\mathcal{G}}(\iota_{0}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\bigl{(}\psi(a)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\bigr{)}+(\iota_{r-1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)$
$=\delta_{\scriptscriptstyle\mathcal{G}}\bigl{(}\iota_{0}(\eta_{\mathcal{G}}^{*}a)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\bigr{)}+(\iota_{r-1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)$
$=\Phi_{r}\bigl{(}\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}(\eta_{\mathcal{G}}^{*}a)+\iota_{r-1}(a)\bigr{)}=\Phi_{r}\bigl{(}A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{r}(a)\bigr{)}.$
Finally, we note that if $\{e_{1},\dotsc,e_{\dim{\operatorname{C}(\mathcal{G})}}\}$ is an orthonormal basis of $\operatorname{C}(\mathcal{G})$ then for any $k\in\{1,\dotsc,r\}$
$\Phi_{r}\bigl{(}\iota_{k}(e_{j})\bigr{)}=(\iota_{k}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(e_{j})=\sum_{i=1}^{r}\iota_{k}(e_{i})\otimes u_{i,j}$
showing that $\bigl{\{}(\omega\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\Phi_{r}(x)\,\bigr{|}\bigl{.}\,x\in\operatorname{C}(\mu_{r-1}(\mathcal{G})),\>\omega\in\operatorname{C}(\mu_{r-1}(\mathcal{G}))^{*}\bigr{\}}$ generates $\mathrm{C}^{*}$-algebra $\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$. ∎
We can amplify actions of quantum groups on $\operatorname{C}(\mathcal{G})$ into those on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ for a Hilbert space $\mathscr{H}$. In fact, the action $\rho_{\scriptscriptstyle\mathcal{G}}$ of $\operatorname{QAut}(\mathcal{G})$ on $\mathcal{G}$ defines
$\rho_{\mathcal{G},\mathscr{H}}=(\mathrm{id}_{\operatorname{C}(\mathcal{G})}\otimes\boldsymbol{\chi})(\rho_{\mathcal{G}}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})})\colon\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\longrightarrow\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G})),$
the action of $\operatorname{QAut}(\mathcal{G})$ on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$. For a partition of unity $\mathfrak{P}\subset\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$, we consider the action $\bigl{(}\rho_{\mathcal{G},\mathscr{H}}\bigr{)}_{\mathfrak{P}}$ of $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$.
Lemma 3.2.
Let $\mathfrak{P}=\{P_{1},\ldots,P_{n}\}$ be a partition of unity in $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$, and take $r\geq 2$. Then
$Q_{0}$ $=\iota_{0}\iota^{*}_{0}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})},$
$Q_{a}$ $=\sum\limits_{k=1}^{r}(\iota_{k}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})})P_{a}(\iota_{k}^{*}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}),\quad a=1,\dotsc,n.$
form a partition of unity $\mathfrak{P}_{\mu}^{r-1}$ in $\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})$.
Proof.
First, notice that $Q_{0}$ and $Q_{a}$ with $a=1,\ldots,n$ are projections. Indeed, we immediately notice that both $Q_{0}$ and $Q_{a}$ with $a=1,\ldots,n$ are idempotent. Furthermore, since $\iota_{k}$, $k=0,\ldots,r$ are $\ast$-maps, we see that all elements of $\mathfrak{P}_{\mu}^{r-1}$ are self-adjoint. Finally, we note that
$\sum\limits_{Q\in\mathfrak{P}_{\mu}^{r-1}}\!\!\!\!Q$ $=Q_{0}+\sum\limits_{a=1}^{n}Q_{a}^{k}=\bigl{(}\iota_{0}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr{)}\bigl{(}\iota_{0}^{\ast}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})})+\sum\limits_{k=1}^{r}\sum\limits_{a=1}^{n}\bigl{(}\iota_{k}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr{)}P_{a}(\iota_{k}^{\ast}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})})$
$=\Bigl{(}\iota_{0}\iota_{0}^{\ast}+\sum\limits_{k=1}^{r}\iota_{k}\iota_{k}^{\ast}\Bigr{)}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}=\mathds{1}_{\operatorname{C}\bigl{(}\mu_{r-1}(G)\bigr{)}\otimes\operatorname{B}(\mathscr{H})}.$
We refer to $\mathfrak{P}_{\mu}^{r-1}$ as the induced partition of unity. For $r=2$, we write $\mathfrak{P}_{\mu}\equiv\mathfrak{P}_{\mu}^{1}$. If $\mathscr{H}$ is one-dimensional, we deal with partitions of unity in the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathcal{G})$.
Corollary 3.3.
Let $\mathcal{G}$ be a quantum graph and $\mathfrak{P}=\{P_{1},\ldots,P_{n}\}\subset\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ be a partition of unity. Then for any $r\geq{2}$, the compact quantum group $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ is embedded as a closed quantum subgroup of $\operatorname{QAut}(\mu_{r-1}(\mathcal{G}))_{\mathfrak{P}_{\mu}^{r-1}}$.
Proof.
Let $\Phi_{r,\mathscr{H}}$ be the homomorphism constructed in Proposition 3.1 extended onto $\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})$, that is for $x\in\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})$ we define
$\Phi_{r,\mathscr{H}}(x)=\bigl{(}\iota_{0}\iota_{0}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}(x)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}+\sum\limits_{k=1}^{r}\bigl{(}\iota_{k}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\bigr{)}\rho_{\mathcal{G},\mathscr{H}}\bigl{(}\iota_{k}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}(x).$
Further, we note that
$\Phi_{r,\mathscr{H}}(Q_{0})=\bigl{(}\iota_{0}\iota_{0}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}\bigl{(}\iota_{0}\iota_{0}^{\ast}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr{)}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}=Q_{0}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))},$
and
$\Phi_{r,\mathscr{H}}(Q_{a})$ $=\sum\limits_{k,l=1}^{r}\bigl{(}\iota_{k}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\bigr{)}\rho_{\mathcal{G},\mathscr{H}}\bigl{(}\iota_{k}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}\bigl{(}\iota_{l}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr{)}P_{a}\bigl{(}\iota_{l}^{\ast}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr{)}$
$=\sum\limits_{k=1}^{r}\bigl{(}\iota_{k}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\bigr{)}\rho_{\mathcal{G},\mathscr{H}}(P_{a})(\iota_{k}^{\ast}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}).$
If $\rho_{\mathcal{G},\mathscr{H}}(P_{a})=P_{a}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$, then the above expression simplifies into $Q_{a}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$, so that $\Phi_{r,\mathscr{H}}$ induces a homomorphism
$\Phi_{r,\mathscr{H}}^{\mathfrak{P}}\colon\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})\longrightarrow\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}).$
By the universal property of $\operatorname{QAut}(\mu_{r-1}(\mathcal{G}))$ there exists a unique homomorphism
$\pi_{r}^{\mathfrak{P}}\colon\operatorname{C}(\operatorname{QAut}(\mu_{r-1}(\mathcal{G})))\longrightarrow\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}})$
such that $\Phi_{r,\mathscr{H}}^{\mathfrak{P}}=(\mathrm{id}\otimes\pi_{r}^{\mathfrak{P}})\circ\rho_{\scriptscriptstyle\mu_{r-1}(\mathcal{G}),\mathscr{H}}$. Moreover, $\pi_{r}^{\mathfrak{P}}$ restricts itself to a well-defined homomorphism $\pi_{r}^{\mathfrak{P},\mathfrak{P}_{\mu}^{r-1}}\colon\operatorname{C}(\operatorname{QAut}(\mu_{r-1}(\mathcal{G}))_{\mathfrak{P}_{\mu}^{r-1}})\to\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}})$ and this homomorphism is compatible with comultiplications (cf. [19, Proposition 4.7]) and is surjective by the last statement of Proposition 3.1. ∎
3.1. The role of twin vertices
Definition 3.4.
Let $G=(V,E)$ be a classical graph. For $v\in V$ define $N(v)=\bigl{\{}u\in{V}\,\bigr{|}\bigl{.}\,\{u,v\}\in{E}\bigr{\}}$. Two vertices $x,y\in V$ are said to be twins if $N(x)=N(y)$.
Theorem 3.5 ([1, Proposition 4.1]).
If $G=(V,E)$ is a twin-free finite graph, then every automorphism of $\mu(G)$ that fixes the point $\bullet$ is induced by an automorphism of $G$.
We will present here a slightly modified proof of this theorem which makes more explicit use of the adjacency matrix of $G$, allowing us to extend the reasoning to the case of quantum graphs in Theorem 3.10.
Proof of Theorem 3.5.
Let $\widetilde{\rho}$ be such an automorphism of $\mu(G)$ that fixes $\bullet$. It follows that the matrix of $\widetilde{\rho}$ as a linear map $\mathbb{C}\oplus\operatorname{C}(V)\oplus\operatorname{C}(V)\to\mathbb{C}\oplus\operatorname{C}(V)\oplus\operatorname{C}(V)$ must be of the form
$\widetilde{\rho}=\begin{bmatrix}1&0&0\\ 0&\rho_{11}&\rho_{12}\\ 0&\rho_{21}&\rho_{22}\end{bmatrix}.$
From the commutation with the adjacency matrix,
$A_{\mu(G)}=\begin{bmatrix}0&0&1\\ 0&A_{G}&A_{G}\\ 1&A_{G}&0\end{bmatrix}$
we deduce that $\rho_{12}=\rho_{21}=0$. Denoting $\rho_{i}=\rho_{ii}\in\operatorname{Aut}(G)$, for $i=1,2$, the remaining conditions are
$\rho_{2}A_{G}=A_{G}\rho_{1}=\rho_{1}A_{G}=A_{G}\rho_{2}.$ (3)
Now, evaluating both sides of the equation $\rho_{1}A_{G}=\rho_{2}A_{G}$ on $\delta_{v}$, the characteristic function of the vertex $v\in V$, we end up with equation $\rho_{1}\chi_{N(v)}=\rho_{2}\chi_{N(v)}$, where $\chi_{C}$ is the indicator function of the set $C$. Since $\rho_{1},\rho_{2}$ are automorphisms of the classical graph $G$, there exist permutations $\tau,\sigma$ of $V$ such that $\rho_{1}\chi_{N(v)}=\chi_{N(\tau(v))}$ and $\rho_{2}\chi_{N(v)}=\chi_{N(\sigma(v))}$. The lack of twin vertices implies that $\sigma=\tau$, so that $\rho_{1}=\rho_{2}$, i.e. every automorphism of $\mu(G)$ that has $\bullet$ as a fixed point is induced by an automorphism of $G$. ∎
Remark 3.6.
Note that in the proof presented above the conclusion $\rho_{1}=\rho_{2}$ can be immediately reached from (3) if we assume that $A_{G}$ is invertible. This is of course a stronger assumption than $G$ being twin-free.
Remark 3.7.
A graph $G=(V,E)$ has twin vertices if and only if its adjacency matrix $A_{G}$ has at least two same rows or, equivalently, columns. The existence of such two rows is equivalent to the existence of a permutation $\pi$ of the set $V$ such that $P_{\pi}A_{G}=A_{G}$, where $P_{\pi}$ is the corresponding permutation matrix. In contrast, the presence of the same columns is equivalent to the existence of $\sigma$ such that $A_{G}P_{\sigma}=A_{G}$.
Motivated by the above observation, we introduce the following definition:
Definition 3.8.
Let $\mathcal{G}=(\mathbb{V}_{\mathcal{G}},\psi_{\mathcal{G}},A_{\scriptscriptstyle\mathcal{G}})$ be a quantum graph and $\mathbb{H}$ be a compact quantum group. We say that $\mathcal{G}$ has $\mathbb{H}$-twin vertices, or simply $\mathbb{H}$-twins, if there exists a pair of different actions of $\mathbb{H}$ on $\mathcal{G}$
$\rho_{1},\rho_{2}\colon\operatorname{C}(\mathcal{G})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{C}(\mathbb{H})$
such that
$\rho_{1}\circ A_{\scriptscriptstyle\mathcal{G}}=\rho_{2}\circ A_{\scriptscriptstyle\mathcal{G}}.$ (4)
We say that $\mathcal{G}$ has quantum twin vertices if there exists a quantum group $\mathbb{H}$ such that $\mathcal{G}$ has $\mathbb{H}$-twins. If a classical group $H$ that meets the above conditions exists, then we say that $\mathcal{G}$ has classical twin vertices.
Remark 3.9.
Suppose that $\mathcal{G}$ and $\mathbb{H}$ are classical, i.e. $\mathbb{H}$ can be identified with a compact group $H$. Suppose that $\mathcal{G}$ has $H$-twins, i.e. there exist $\rho_{1},\rho_{2}\colon\operatorname{C}(\mathcal{G})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{C}(\mathbb{H})$ satisfying (4). The actions $\rho_{1}$ and $\rho_{2}$ can be viewed as homomorphisms $H\ni h\mapsto\rho_{i}(h)\in\operatorname{Aut}(\mathcal{G})$. Since $\rho_{1}\neq\rho_{2}$ there is $h\in H$ such that $\rho_{1}(h)\neq\rho_{2}(h)$ and
$\rho_{1}(h)\circ A_{\scriptscriptstyle\mathcal{G}}=\rho_{2}(h)\circ A_{\scriptscriptstyle\mathcal{G}}.$
Defining $\sigma=\rho_{2}(h)^{-1}\rho_{1}(h)\in\operatorname{Aut}(\mathcal{G})$ we have $\sigma\circ A_{\scriptscriptstyle\mathcal{G}}=A_{\scriptscriptstyle\mathcal{G}}$. Since $\sigma\neq\mathrm{id}$ we can find two vertices $v_{1}\neq v_{2}$ and $\sigma(v_{1})=v_{2}$ and it is easy to see that $v_{1}$ and $v_{2}$ are twins.
Conversely, if $v_{1},v_{2}\in\mathbb{V}_{\mathcal{G}}$ are twins we can define an order two symmetry $\pi\in\operatorname{Aut}(\mathcal{G})$ such that $\pi(v)=v$ for all $v\in\mathbb{V}\setminus\{v_{1},v_{2}\}$ and $\pi(v_{1})=v_{2}$ (note that this is a different way of saying that $v_{1}$ and $v_{2}$ are twin vertices). Taking $\mathbb{H}=\mathbb{Z}_{2}=\{e,h\}$ and putting $\rho_{1}(h)=\pi$ and $\rho_{2}(h)=\mathrm{id}$ we see that $\mathcal{G}$ has $\mathbb{Z}_{2}$-twins in the sens of Definition 3.8.
Next, we generalize Theorem 3.5 to the case of quantum graphs.
Theorem 3.10.
Let $\mathcal{G}=(\mathbb{V}_{\mathcal{G}},\psi_{\mathcal{G}},A_{\scriptscriptstyle\mathcal{G}})$ be a quantum graph that does not have quantum twin vertices. Let $\mathfrak{P}$ be a partition of unity in $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ with some (separable) Hilbert space $\mathscr{H}$. Then $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}\cong\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$.
Proof.
By Corollary 3.3, $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ is a closed quantum subgroup of $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}$. Denote the corresponding quotient map by $\pi_{2}^{\mathfrak{P},\mathfrak{P}_{\mu}}$. On the other hand, let
$\phi\colon\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})\longrightarrow\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}\bigl{(}\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}\bigr{)}$
be the action of $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}$ on $\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})$.
Then $\phi=(\mathrm{id}\otimes\pi_{\mu})\rho_{\mu(\mathcal{G}),\mathscr{H}}$ with the quotient map $\pi_{\mu}\colon\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G})))\to\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}})$. For $i,j\in\{0,1,2\}$, define $\phi_{i,j}\colon\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}})$ by
$\phi_{i,j}=(\iota_{i}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mu\circ(\mathcal{G}))_{\mathfrak{P}_{\mu}})})\circ\phi\circ\bigl{(}\iota_{j}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}.$
It is easy to see that since $\left[\begin{smallmatrix}1\\ 0\\ 0\end{smallmatrix}\right]$ belongs to $\mathfrak{P}_{\mu}$ and $\left[\begin{smallmatrix}0\\ \mathds{1}\\ \mathds{1}\end{smallmatrix}\right]$ is a sum of elements of $\mathfrak{P}_{\mu}$, we have $\phi_{0,k}=0$ and $\phi_{k,0}=0$ for $k=1,2$, and $\phi_{0,0}(1)=1\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$. For $k,l=1,2$ let
$\widetilde{\phi}_{k,l}=(\mathrm{id}_{\operatorname{C}(\mathcal{G})}\otimes\pi_{2}^{\mathfrak{P},\mathfrak{P}_{\mu}})\phi_{k,l}\colon\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}).$
Then
$\widetilde{\phi}_{k,l}=\bigl{(}\iota_{k}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\Pi\bigr{)}\rho_{\mu(\mathcal{G})}\bigl{(}\iota_{l}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)},$
with $\Pi=\pi_{2}^{\mathfrak{P},\mathfrak{P}_{\mu}}\circ\pi_{\mu}$. Recall that
$A_{\mu(\mathcal{G})}=\delta_{\mathcal{G}}\iota_{2}\eta_{\mathcal{G}}\iota_{0}^{\ast}+\delta_{\mathcal{G}}\iota_{0}\eta_{\mathcal{G}}^{\ast}\iota_{2}^{\ast}+\iota_{1}A_{\mathcal{G}}\iota_{1}^{\ast}+\iota_{1}A_{\mathcal{G}}\iota_{2}^{\ast}+\iota_{2}A_{\mathcal{G}}\iota_{1}^{\ast},$
and $\rho_{\mu(\mathcal{G})}A_{\mu(\mathcal{G})}=\bigl{(}A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G})))}\bigr{)}\rho_{\mu(\mathcal{G})}$ which leads to
$\rho_{\mu(\mathcal{G}),\mathscr{H}}\bigl{(}A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}=\bigl{(}A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}\bigr{)}\rho_{\mu(\mathcal{G}),\mathscr{H}}.$
Next, we explore the equation
$\bigl{(}\iota_{k}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\Pi\bigr{)}$ $\rho_{\mu(\mathcal{G}),\mathscr{H}}(A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})})\bigl{(}\iota_{l}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}$
$=(\iota_{k}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\Pi)\bigl{(}A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G})))}\bigr{)}\rho_{\mu(\mathcal{G}),\mathscr{H}}\bigl{(}\iota_{l}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}$
for $k,l=0,1,2$. First, taking $(k,l)=(0,1)$ and $(k,l)=(1,0)$ leads to $\widetilde{\phi}_{2,1}=\widetilde{\phi}_{1,2}=0$. Using this and considering $(k,l)=(1,2)$ gives $\widetilde{\phi}_{1,1}A_{\mathcal{G}}=A_{\mathcal{G}}\widetilde{\phi}_{2,2}$, while $(k,l)=(2,1)$ leads to $\widetilde{\phi}_{2,2}A_{\mathcal{G}}=A_{\mathcal{G}}\widetilde{\phi}_{1,1}$. Finally, $(k,l)=(1,1)$ leads to $\widetilde{\phi}_{1,1}A_{\mathcal{G}}=A_{\mathcal{G}}\widetilde{\phi}_{1,1}$. Therefore, $\widetilde{\phi}_{1,1}A_{\mathcal{G}}=\widetilde{\phi}_{2,2}A_{\mathcal{G}}$. Since $\mathcal{G}$ does not have quantum twin vertices, we deduce $\widetilde{\phi}_{1,1}=\widetilde{\phi}_{2,2}$. This ends the proof. ∎
3.2. Quantum twin vertices in small classical graphs
It is unclear whether the notion of quantum twin vertices is broader than the usual concept of twin vertices when applied to classical graphs. In this section, we consider this question for graphs with few vertices. In view of Remark 3.6 we will be mainly interested in graphs with no twins, but whose adjacency matrix is not invertible.
Considering graphs up to the classical isomorphism, we begin with the following observation.
Proposition 3.11.
There is no graph $G$ with $|V(G)|\leq 5$ such that $G$ has quantum symmetries, has no twins, and $\det A_{G}=0$.
Proof.
The statement is clear for $|V(G)|\leq 3$. For $|V(G)|=4$, only three graphs possess quantum symmetries, cf. [15]. Their adjacency matrices are
$\begin{bmatrix}0&0&1&1\\ 0&0&1&1\\ 1&1&0&0\\ 1&1&0&0\end{bmatrix},\begin{bmatrix}0&0&1&1\\ 0&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{bmatrix},\begin{bmatrix}0&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{bmatrix}$
respectively. The first two have twins, as the corresponding adjacency matrices have two identical rows. However, the last one is twin-free, but the adjacency matrix has non-zero determinant.
For $|V(G)|=5$, only ten graphs possess quantum symmetries (again see [15]). One of them is the complete graph $K_{5}$, and the adjacency matrices of the remaining ones are
$A_{G_{1}^{5}}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{G_{2}^{5}}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&1&1\\ 0&0&1&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{G_{3}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&0&1&1\\ 0&0&0&1&1\\ 1&1&1&0&0\\ 1&1&1&0&0\\ \end{bmatrix},$
$A_{G_{4}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&0&1&1\\ 0&0&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{G_{5}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&0&1\\ 0&1&0&0&1\\ 1&0&0&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{G_{6}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&1&1\\ 0&1&0&1&1\\ 1&1&1&0&0\\ 1&1&1&0&0\end{bmatrix},$
$A_{G^{5}_{7}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&1&1\\ 0&1&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{G^{5}_{8}}=\begin{bmatrix}0&0&1&1&1\\ 0&0&1&1&1\\ 1&1&0&0&1\\ 1&1&0&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{G^{5}_{9}}=\begin{bmatrix}0&0&1&1&1\\ 0&0&1&1&1\\ 1&1&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\end{bmatrix}.$
Among them, only $K_{5}$, $G_{5}^{5}$ and $G_{7}^{5}$ have no twins. However, we have $\det A_{K_{5}}=4$, $\det A_{G_{5}^{5}}=-4$ and $\det A_{G_{7}^{5}}=-2$. ∎
For graphs with six vertices, we have the following proposition:
Proposition 3.12.
For graphs with $|V(G)|\leq 6$, the only one (up to classical isomorphism) that has both quantum symmetries, no twins, and such that $\det A_{G}=0$ is described by the following adjacency matrix
$A_{G_{6}}=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&1\\ 1&0&0&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}.$
We present the proof of Proposition 3.12 in the Appendix. Here we notice that this graph has the property that for every pair of vertices sharing the same number of edges, there is an edge between them. For example, vertices 2 and 3 have three edges, and there is an edge connecting these two vertices. This property ensures that there are no classical twins.
Proposition 3.13.
The graph $G_{6}$ from Proposition 3.12 does not possess quantum twins.
Proof.
First, by the Fulton criterion ([12], see also [15, Lemma 2.6]) and the fact that for all $1\leq i\leq 6$ we have $\sum\limits_{k=1}^{6}u_{ik}=\sum\limits_{k=1}^{6}u_{ki}=1$, we know that the generating matrix $u$ of $\operatorname{C}(\operatorname{QAut}(G_{6}))$ is of the form
$u=\begin{bmatrix}u_{11}&0&0&0&1-u_{11}&0\\ 0&u_{22}&1-u_{22}&0&0&0\\ 0&1-u_{22}&u_{22}&0&0&0\\ 0&0&0&u_{33}&0&1-u_{33}\\ 1-u_{11}&0&0&0&u_{11}&0\\ 0&0&0&1-u_{33}&0&u_{33}\end{bmatrix}.$
Suppose there exists $\rho_{1},\rho_{2}$, a pair of different actions of a certain quantum group $\mathbb{H}$ on $G_{6}$ such that $\rho_{1}\circ A_{G_{6}}=\rho_{2}\circ A_{G_{6}}$. Let $P$ and $Q$ be matrix representations of $\rho_{1}$ and $\rho_{2}$ on the standard basis of $L^{2}(G)$, respectively. They can be parameterized as
$P=\begin{bmatrix}p_{1}&0&0&0&1-p_{1}&0\\ 0&p_{2}&1-p_{2}&0&0&0\\ 0&1-p_{2}&p_{2}&0&0&0\\ 0&0&0&p_{3}&0&1-p_{3}\\ 1-p_{1}&0&0&0&p_{1}&0\\ 0&0&0&1-p_{3}&0&p_{3}\end{bmatrix}$
and
$Q=\begin{bmatrix}q_{1}&0&0&0&1-q_{1}&0\\ 0&q_{2}&1-q_{2}&0&0&0\\ 0&1-q_{2}&q_{2}&0&0&0\\ 0&0&0&q_{3}&0&1-q_{3}\\ 1-q_{1}&0&0&0&q_{1}&0\\ 0&0&0&1-q_{3}&0&q_{3}\end{bmatrix},$
respectively, where $p_{1},p_{2},p_{3}$ and $q_{1},q_{2},q_{3}$ are orthogonal projections in $\operatorname{C}(\mathbb{H})$. Since they must commute with the adjacency matrix, we get $p_{3}=p_{1}$ and $q_{3}=q_{1}$. With this condition, the relation $\rho_{1}\circ A_{G_{6}}=\rho_{2}\circ A_{G_{6}}$ leads to $p_{1}=q_{1}$ and $p_{2}=q_{2}$, so $\rho_{1}$ is identical to $\rho_{2}$, a contradiction. ∎
Corollary 3.14.
The notions of quantum and classical twin vertices coincide for graphs $G$ with $|V(G)|\leq 6$.
We now consider a subclass of classical graphs, the vertex-transitive graphs. We recall that a graph $G$ is called vertex-transitive if $\operatorname{Aut}(G)$ acts transitively on the set of its vertices. It is known[3, 8, 9, 17, 18] exactly which of these graphs with $|V|\leq 13$ possess nonclassical quantum symmetries. Using these results, we have the following:
Proposition 3.15.
For connected vertex-transitive graphs with at most $13$ vertices, only $K_{2}\Box K_{5}$ and $C_{4}\Box C_{3}$ possess non-classical quantum symmetries, have no twin vertices and vanishing determinant of the adjacency matrix.
Proof.
We analyze the properties of such graphs with non-classical quantum symmetries, listed in [17, Tables 1-2]. We first observe that among them, the only ones with the vanishing determinant of the adjacency matrix are $K_{2}\Box K_{5}$, $C_{10}(4)$, $C_{2}\Box C_{4}$ and $C_{12}(5)$. However, $C_{10}(4)$ possesses five pairs of twin vertices, while $C_{12}(5)$ has six such pairs. For the remaining two graphs, we have
$A_{K_{2}\Box K_{5}}=A_{K_{2}}\otimes\mathds{1}_{5}+\mathds{1}_{2}\otimes A_{K_{5}}=\begin{bmatrix}0&1&1&1&1&1&0&0&0&0\\ 1&0&1&1&1&0&1&0&0&0\\ 1&1&0&1&1&0&0&1&0&0\\ 1&1&1&0&1&0&0&0&1&0\\ 1&1&1&1&0&0&0&0&0&1\\ 1&0&0&0&0&0&1&1&1&1\\ 0&1&0&0&0&1&0&1&1&1\\ 0&0&1&0&0&1&1&0&1&1\\ 0&0&0&1&0&1&1&1&0&1\\ 0&0&0&0&1&1&1&1&1&0\end{bmatrix}$
and
$A_{C_{4}\Box C_{3}}=A_{C_{4}}\otimes\mathds{1}_{3}+\mathds{1}_{4}\otimes A_{C_{3}}=\left[\begin{array}[]{cccccccccccc}0&1&1&1&0&0&0&0&0&1&0&0\\ 1&0&1&0&1&0&0&0&0&0&1&0\\ 1&1&0&0&0&1&0&0&0&0&0&1\\ 1&0&0&0&1&1&1&0&0&0&0&0\\ 0&1&0&1&0&1&0&1&0&0&0&0\\ 0&0&1&1&1&0&0&0&1&0&0&0\\ 0&0&0&1&0&0&0&1&1&1&0&0\\ 0&0&0&0&1&0&1&0&1&0&1&0\\ 0&0&0&0&0&1&1&1&0&0&0&1\\ 1&0&0&0&0&0&1&0&0&0&1&1\\ 0&1&0&0&0&0&0&1&0&1&0&1\\ 0&0&1&0&0&0&0&0&1&1&1&0\\ \end{array}\right].$
By a direct inspection, we verify that the above matrices do not have two identical rows so there are no classical twin vertices in these graphs. ∎
We note that the Fulton criterion does not yield any information about the graphs $K_{2}\Box K_{5}$ (the so-called prism having a basis in $K_{5}$, $\mathrm{Pr}(K_{5})$ [3, Definition 4.2]) and $C_{4}\Box C_{3}$ as they are vertex-transitive [15, Remark 2.7]. The other potential candidate is the Tesseract graph with $16$ vertices. Its adjacency matrix is singular, while it does not have classical twins. The classical symmetry group is $H_{4}$, while the quantum symmetry group is $O_{4}^{-1}$, see [4].
Question 3.16.
Do the above-mentioned graphs possess quantum twin vertices?
4. The distinguishing number
A labeling of a classical graph $G=(V,E)$ is a mapping $V\to\{1,\dotsc,c\}$, while a coloring is a labeling with the property that adjacent vertices are labeled with different integers (colors). The chromatic number of $G$ defined as the smallest $c$ such that there exists a coloring of $G$ with $c$ colors is an invariant that measures how “well connected” the graph $G$ is. A different invariant called distinguishing number measures this property with respect to the size of the automorphism group of $G$. We recall here this classical definition:
Definition 4.1 ([1]).
Let $G=(V,E)$ be a finite graph. A $c$-distinguishing labeling of $G$ is a function $\phi\colon V\to\{1,\dotsc,c\}$ such that for every $\sigma\in\operatorname{Aut}(G)\setminus\{\mathrm{id}\}$, there exists $v\in V$ such that $\phi(v)\neq\phi(\sigma(v))$.
The behavior of the distinguishing number and variations of it under the Mycielski transformation was widely studied in recent years; see, e.g. [1, 14]. Here we introduce a version of Definition 4.1 for quantum graphs. To that end, we recall that a quantum graph $\mathcal{G}$ defines an operator system $S_{\scriptscriptstyle\mathcal{G}}\subset\operatorname{B}(L^{2}(\mathcal{G}))$ through $S_{\scriptscriptstyle\mathcal{G}}=\bigl{\{}m(A\otimes{X})m^{*}\,\bigr{|}\bigl{.}\,X\in\operatorname{B}(L^{2}(\mathcal{G}))\bigr{\}}$ (see, e.g. [13, Section 2.4]).
Definition 4.2.
Let $\mathcal{G}$ be an irreflexive quantum graph, $S_{\scriptscriptstyle\mathcal{G}}$ the operator system associated with $\mathcal{G}$ and $\mathscr{H}$ be a Hilbert space. A partition of unity $\{P_{1},\dotsc,P_{c}\}$ in $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ is called a quantum $c$-labeling of $\mathcal{G}$. We say that that a quantum $c$-labeling of $\mathcal{G}$ is a quantum $c$-coloring of $\mathcal{G}$ if
$\forall\>a\in\{1,\dotsc,c\}\quad\forall\>X\in S_{\scriptscriptstyle\mathcal{G}}\quad P_{a}(X\otimes\mathds{1}_{\mathcal{\mathscr{H}}})P_{a}=0.$
Definition 4.3.
Let $\mathfrak{P}=\{P_{1},\dotsc,P_{c}\}\subset\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ be a quantum $c$-labeling of $\mathcal{G}$. We say that $\mathfrak{P}$ is a quantum $c$-distinguishing labeling of $\mathcal{G}$ if the quantum group $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ is trivial.
Remark 4.4.
Note that if $\mathcal{G}$ is a classical graph (i.e. the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathcal{G})$ is commutative) and $\mathfrak{P}$ is a classical $c$-labeling (i.e. $\mathscr{H}$ is one dimensional) then it corresponds to labeling of vertices $f\colon{V_{\mathcal{G}}}\to\{1,\dotsc,c\}$ uniquely defined by the condition $P_{i}(v)=1\Leftrightarrow{f(v)}=i$. In this case $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ can be viewed as a quantum group of automorphisms preserving $f$. In particular, the abelianization of $\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}})$ corresponds to the subgroup of $\operatorname{Aut}(\mathcal{G})$ that preserves $f$.
Definition 4.5.
Let $\mathcal{G}$ be a quantum graph. We define the quantum distinguishing number $D(\mathcal{G})$ of $\mathcal{G}$ as the smallest $c$ such that $\mathcal{G}$ admits a distinguishing $c$-labeling.
Theorem 4.6.
Let $\mathcal{G}$ be a quantum graph with $\dim\operatorname{C}(\mathcal{G})\geq 2$ which does not possess quantum twin vertices. Then $D(\mu(\mathcal{G}))\leq D(\mathcal{G})+1$.
Proof.
Let $\mathfrak{P}=\{P_{1},\dotsc,P_{D(\mathcal{G})}\}\subset\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ be a quantum $D(\mathcal{G})$-distinguishing labeling of $\mathcal{G}$, so that $\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}})=\mathbb{C}$, and consider the induced partition of unity $\mathfrak{P}_{\mu}$ on $\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})$, that is $\mathfrak{P}_{\mu}=\{Q_{0},Q_{1},\ldots,Q_{D(\mathcal{G})}\}$ with
$Q_{0}$ $=\iota_{0}\iota^{*}_{0}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})},$
$Q_{a}$ $=(\iota_{1}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})})P_{a}(\iota_{1}^{*}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})})+(\iota_{2}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})})P_{a}(\iota_{2}^{*}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}),\quad a=1,\dotsc,D(\mathcal{G}).$
By Theorem 3.10, $\mathfrak{P}_{\mu}$ is a quantum $(D(\mathcal{G})+1)$-labeling of $\mu(\mathcal{G})$. ∎
Acknowledgments
The work of AB was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2111 – 390814868. This research was partially supported by the University of Warsaw Thematic Research Program “Quantum Symmetries”. AB is supported by the Alexander von Humboldt Foundation. PS was supported by the NCN (National Science Center, Poland) grant no. 2022/47/B/ST1/00582.
Furthermore, the authors wish to thank Adam Skalski for the fruitful discussion on the subject of quantum symmetries of quantum graphs.
Appendix: Proof of Proposition 3.12
Here we present the proof of Proposition 3.12 explicitly listing all $55$ graphs with six vertices that have quantum symmetries ([15]), and for each of them verifying if it possesses twins. In case the graph is twin-free, we provide instead the value of the determinant of the corresponding adjacency matrix:
$A_{G_{1}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{2}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{3}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&1&1&0&1\\ 1&1&0&0&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{4}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{5}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{6}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&1&0&1&0\\ 0&0&1&1&0&1\\ 1&1&0&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{7}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&1&0&1&1\\ 0&0&1&1&0&1\\ 1&1&0&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{8}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&0&1&1&0&0\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{9}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&1&1&1&0&0\\ 1&0&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{10}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&0&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{11}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&1&1&0&0\\ 1&0&1&1&0&0\end{bmatrix},\quad\det=-1,$ $A_{G_{12}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&1&1&1&0\end{bmatrix},\quad\det=-1,$
$A_{G_{13}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 0&1&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=-1,$ $A_{G_{14}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&0&1&1&0\\ 0&1&1&0&0&1\\ 0&1&1&0&0&1\\ 1&0&0&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{15}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&0&1&1&0\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&0&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{16}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&1\\ 0&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{17}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&0\\ 0&1&1&1&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{18}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{19}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&1&1&1&0\\ 0&1&0&1&1&0\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=3,$ $A_{G_{20}^{6}}$ $=\begin{bmatrix}0&0&0&0&0&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 0&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$
$A_{G_{21}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{22}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{23}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{24}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{25}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ 1&1&0&0&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{26}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&0&1\\ 0&0&1&0&1&0\\ 1&1&0&1&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{27}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{28}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{29}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&0\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{30}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{31}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{32}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{33}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&0&0\\ 0&1&1&0&0&1\\ 1&0&0&0&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=3,$ $A_{G_{34}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&0&0\\ 0&1&1&0&1&1\\ 1&0&0&1&0&1\\ 1&0&0&1&1&0\end{bmatrix},\quad\det=4,$
$A_{G_{35}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&0\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&0&1&1&0&1\\ 1&0&1&1&1&0\end{bmatrix},\quad\det=4,$ $A_{G_{36}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&0&1\\ 1&0&0&0&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=7,$
$A_{G_{37}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&0\\ 1&0&0&1&0&0\\ 1&1&1&0&0&0\end{bmatrix},\quad\det=-5,$ $A_{G_{38}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&1\\ 1&0&0&1&0&0\\ 1&1&1&1&0&0\end{bmatrix},\quad\det=0,$
$A_{G_{39}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{40}^{6}}$ $=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 0&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=4,$
$A_{G_{41}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\\ 1&1&1&0&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{42}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{43}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&0&1\\ 1&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{44}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 0&0&0&1&1&1\\ 1&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$\ A_{G_{45}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&0&1&1\\ 1&0&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{46}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&0&1&1\\ 1&0&0&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=7,$
$A_{G_{47}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&1&1&1\\ 1&0&1&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{48}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&0&1&1\\ 0&1&0&1&1&1\\ 1&0&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$
$A_{G_{49}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&0&0\\ 1&1&1&0&0&1\\ 1&1&1&0&1&0\end{bmatrix},\quad\det=-1,$ $A_{G_{50}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&0&1\\ 1&1&1&0&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{51}^{6}}$ $=\begin{bmatrix}0&0&0&1&1&1\\ 0&0&1&1&1&1\\ 0&1&0&1&1&1\\ 1&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix},\quad\det=3,$ $A_{G_{52}^{6}}$ $=\begin{bmatrix}0&0&1&1&1&1\\ 0&0&1&1&1&1\\ 1&1&0&0&1&1\\ 1&1&0&0&1&1\\ 1&1&1&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}\quad\text{has twins},$
$A_{G_{53}^{6}}$ $=\begin{bmatrix}0&0&1&1&1&1\\ 0&0&1&1&1&1\\ 1&1&0&0&1&1\\ 1&1&0&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$ $A_{G_{54}^{6}}$ $=\begin{bmatrix}0&0&1&1&1&1\\ 0&0&1&1&1&1\\ 1&1&0&1&1&1\\ 1&1&1&0&1&1\\ 1&1&1&1&0&1\\ 1&1&1&1&1&0\end{bmatrix}\quad\text{has twins},$
and $G_{55}^{6}=K_{5}$ with $\det K_{5}=-5$. The above list shows that only $G_{38}^{6}$ possesses all the required properties.